# A Survey on SAR ship classification using Deep Learning

### v1

@@ VII-A Addressing Data scarcity
The data used by researchers for DL-based SAR ship classification is seldom made public due to privacy or digital rights reasons. Future research should focus on developing a methodology to publicize the data without any such problem. Also, developing more robust data augmentation techniques can result in diverse SAR datasets. Collaborative efforts of standardizing and sharing the data can significantly benefit the research community.
@@ VII-C Handcrafted and Deep Features Integration
DL models have achieved remarkable success in feature learning for various tasks. However, handcrafted features, meticulously designed to capture specific characteristics of SAR images, remain a valuable tool. Further research is warranted to explore how these can be effectively integrated with DL models for SAR image classification. This combined approach has the potential to leverage the strengths of both machine learning and DL methods, potentially leading to significant performance improvements.
@@ VII-D Standardized Performance Metrics
The field of SAR ship classification currently lacks a standardized set of performance metrics, hindering comprehensive comparisons between different studies. Utilizing metrics like F1-score can partially address this issue, but future research should focus on establishing a common set of metrics specifically tailored to the nuances of SAR ship classification tasks. This would enable more robust evaluations and facilitate meaningful comparisons between different deep learning models and training strategies.
@@ VII-E Interpretability of DL models
While interpretability has not always been a central focus in SAR ship classification research, it presents a valuable avenue for future exploration. By leveraging interpretability techniques, researchers can gain deeper insights into the features most critical for the model’s decisions. This enhanced understanding can lead to more informed analysis, improved trust in model predictions, and potentially even the identification of potential biases within the model.
@@ VII-F Interdisciplinary Collaborations
By focusing on these directions, future research can build on our findings and continue to advance the field of SAR ship classification using deep learning, ultimately improving the accuracy and efficiency of maritime surveillance and related applications.
@@ VIII Conclusion
This survey has systematically reviewed the diverse techniques employed to train deep learning models for SAR ship classification. We identify critical trends and challenges in this domain, emphasizing the importance of integrating handcrafted features, utilizing public datasets, data augmentation, fine-tuning, explainability, and interdisciplinary collaborations for enhancing DL model performance.
We established a survey methodology with clearly defined research questions. Our literature search yielded a comprehensive dataset of relevant papers, which were then categorized into a first-of-its-kind taxonomy encompassing deep learning techniques, handcrafted feature use, SAR attribute utilization, and the impact of fine-tuning.
Section IV provides detailed explanations of the included papers, categorized into the four sections (discussing DL techniques, impact of handcrafted features, use of SAR properties and fine-tuning). Section V presents the methodologies employed in SAR ship classification. Finally, Section VII outlines potential avenues for addressing current challenges, such as data scarcity, novel DL architectures, interpretability, performance metrics, feature integration, and interdisciplinary collaborations.

### v2

@@ VII-A Addressing Data scarcity
The data used by researchers for DL-based SAR ship classification is seldom made public due to privacy or digital rights reasons. Future research should focus on developing methodologies to publicize the data without such constraints. In addition, creating high-quality datasets from open-access SAR imagery and establishing methodologies for reliable AIS–SAR mapping can play a key role in alleviating the problem of data scarcity. Developing more robust data augmentation techniques can further contribute to the diversity of SAR datasets [72]. Finally, collaborative efforts aimed at standardizing and sharing datasets would significantly benefit the research community.
@@ VII-C Handcrafted and Deep Features Integration
DL models have achieved remarkable success in feature learning for various tasks. However, handcrafted features, meticulously designed to capture specific characteristics of SAR images, remain a valuable tool. Further research is warranted to explore how these can be effectively integrated with DL models for SAR image classification. This combined approach has the potential to leverage the strengths of both machine learning and DL methods, potentially leading to significant performance improvements. Moreover, leveraging handcrafted features and intrinsic SAR properties may provide promising pathways for the development of next-generation models, including foundation models and vision–language architectures [110].
@@ VII-D Standardized Performance Metrics
The field of SAR ship classification currently lacks a standardized set of performance metrics, hindering comprehensive comparisons between different studies. Utilizing metrics like the F1 score can partially address this issue, but future research should focus on establishing a common set of metrics specifically tailored to the nuances of SAR ship classification tasks. This would enable more robust evaluations and facilitate meaningful comparisons between different DL models and training strategies.
@@ VII-E Interpretability of DL models
While interpretability has not always been a central focus in SAR ship classification research, it presents a valuable avenue for future exploration. By leveraging interpretability techniques, researchers can gain deeper insights into the features most critical for the model’s decisions. This enhanced understanding can lead to more informed analysis, improved trust in model predictions, and even to the identification of potential biases within the model.
@@ VII-F Interdisciplinary Collaborations
By focusing on these directions, future research can build on our findings and continue to advance the field of SAR ship classification using DL, ultimately improving the accuracy and efficiency of maritime surveillance and related applications, we also present common challenges and remedies to those challenges in table-VII.
@@ VIII Conclusion
This survey has systematically reviewed the diverse techniques employed to train DL models for SAR ship classification. We identified critical trends and challenges in this domain, emphasizing the importance of integrating handcrafted features, utilizing public datasets, data augmentation, fine-tuning, explainability, and interdisciplinary collaborations for enhancing DL model performance.
We established a survey methodology with clearly defined research questions. Our literature search yielded a comprehensive dataset of relevant papers, which were then categorized into a first-of-its-kind taxonomy encompassing DL techniques, handcrafted feature use, SAR attribute utilization, and the impact of fine-tuning.
In particular, Section IV provides detailed explanations of the included papers, categorized into the four sections (discussing DL techniques, impact of handcrafted features, use of SAR properties and fine-tuning). Section V presents the methodologies employed in SAR ship classification. Finally, Section VII outlines potential avenues for addressing current challenges, such as data scarcity, novel DL architectures, interpretability, performance metrics, feature integration, and interdisciplinary collaborations.